# User Location Disclosure Fails to Deter Overseas Criticism but Amplifies Regional Divisions on Chinese Social Media

### v2

@@ Abstract
We examine the behavioral effects of a user location disclosure policy implemented by Sina Weibo, China’s largest microblogging platform, using a high-frequency dataset of uncensored user engagement—including tens of thousands of comments—on 165 prominent government and media accounts. Exploiting the platform’s abrupt rollout of IP-based location tags on April 28, 2022, we compare user behavior in comment sections before and after the policy change. Although the policy was publicly justified as a measure to curb misinformation and counter foreign influence, we find no decline in participation by overseas users. Instead, it significantly reduced domestic engagement with local issues outside users’ home provinces, particularly among critical comments. Evidence suggests this effect was not driven by generalized fear or concerns about credibility, but by a rise in regionally discriminatory replies that increased the social cost of cross-provincial engagement. Our findings indicate that identity disclosure tools can produce unintended consequences by activating existing social divisions in ways that reinforce state control without direct censorship.
Keywords Information Control  $\cdot$ Group Division  $\cdot$ Social Media  $\cdot$ Location Disclosure
@@ 1 Introduction
In October 2021, the Cyberspace Administration of China released draft regulations requiring platforms to display users’ IP-based locations—by province for domestic users and by country for overseas users. The Chinese government framed the policy as a measure to curb misinformation and foreign interference, but critics viewed it as another step to constrain speech online (Reuters, 2022; The New York Times, 2022). On March 17, 2022, less than a month after the Russian invasion of Ukraine, Sina Weibo, China’s largest microblogging site, began tagging the locations of users posting about the Russo-Ukrainian war, claiming it aimed to minimize misinformation and ensure content authenticity. However, many believe the real intent was to restrict overseas users from sharing information that contradicted the government’s stance (see Supporting Information, SI). At noon on April 28, Weibo abruptly expanded the policy to all posts and comments. Other major platforms quickly followed. By the end of 2022, location disclosure had become a standard feature of Chinese social media.
@@
This paper examines the causal impact of the location disclosure policy implemented on Weibo. Authoritarian regimes use a range of tools to manage online discourse, including censorship (King et al., 2013a; Lorentzen, 2014; Roberts, 2018), surveillance (Esarey and Xiao, 2011; Pan et al., 2023), and orchestrated engagement (Chen and Xu, 2017a, b; Qin et al., 2017). Increasingly, they also rely on platform-enabled identity disclosure—such as real-name registration or location tagging—to steer behavior without direct repression. These tools make identity attributes visible to others and may influence behavior through reputational or social pressures. While prior research suggests that regional cues can heighten incivility and deepen social divisions (Peng, 2021; Chen et al., 2021; Yu and Margolin, 2024; Guo et al., 2023), their causal impact on online participation remains poorly understood. This study helps fill that gap. Beyond the authoritarian context, this study contributes to a broader literature on how platform architecture shapes user behavior. Prior research shows that design features—such as source trustworthiness ratings (Papakyriakopoulos et al., 2022), state-affiliation tags (Liang et al., 2022; Nassetta and Gross, 2020; Bradshaw et al., 2024), and feed-ranking algorithms (Flaxman et al., 2016; Guess et al., 2023)—can influence exposure, sharing, and engagement, even without changing users’ beliefs. Our findings add to this work by showing that simple identity cues, like location tags, can also suppress participation and segment discourse, highlighting the broader behavioral consequences of platform design choices.
Study the effects of user information disclosure faces empirical challenges. First, data accuracy is difficult to ensure in hihgly censored environments like Weibo, where behavioral signals may reflect censorship rather than genuine user response. Prior work shows that 3% of sensitive posts are deleted within 30 minutes and 90% within 24 hours (King et al., 2013b), introducing survivorship bias. Public datasets such as Weiboscope and Weibo-Cov (Weiboscope Project, 2018; Nguyen, 2020) often rely on post-censorship crawling, which overrepresents high-follower accounts, or uses narrow keyword lists. Location data, when available, typically rely on self-reported tags subject to systematic misclassification (Qin et al., 2018). Second, measurement is limited: most studies preserve only original posts and high-level reposts (e.g., Lu et al., 2021; Qin et al., 2024; Meng et al., 2025), omitting comment threads where user reactions to policy often emerge. Information on which users are subject to disclosure—and what is revealed—is typically missing, making it difficult to measure treatment exposure. Third, credible causal inference is rare. Experimental studies are constrained by ethical concerns, while observational approaches rely on platform-generated signals, such as deletion flags or user tags; these signals are usually governed by opaque and inconsistently applied rules, leading to potential confounding. Moreover, most datasets are scraped at low frequency and do not track accounts or posts over time, which limits their value for studying behavioral dynamics.
To overcome these challenges, we implemented a high-frequency data collection strategy that continuously monitored a set of prominent government-affiliated Weibo accounts and their comment sections. Scraping at approximately five-minute intervals allowed us to capture most comments before censorship or removal, reducing survivorship bias and enabling us to track evolving discussions with high temporal resolution. We focused on top comments and their replies—ranked by the platform based on engagement—to observe how public discourse unfolds in visible spaces. The dataset spans the sudden, unannounced rollout of the location disclosure policy on Weibo, providing a natural interruption for before-and-after comparisons. Its panel structure supports within-account comparisons, while repeated snapshots of the comment section capture shifts in user engagement over time.
2 Data and Measurement
We collaborated with industry experts to design a real-time data collection system that monitored 165 prominent Weibo accounts affiliated with the Chinese government. These include official accounts of government offices and state media outlets—such as newspapers, television stations, and major media outlets—identified in a series of reports on institutional influence in 2021.111These reports include multiple editions of Government Weibo Influence Reports (政务微博影响力报告), Sina Government Weibo Reports (新浪政务微博报告), and Weibo Trending Topics Data Reports (微博热搜榜数据报告). See Table LABEL:table:account_list in the SI for a full list of accounts we collected data from. We focused on these accounts because, following the government’s 2013 crackdown on key opinion leaders (Pan et al., 2025), the comment sections of these accounts—due to high traffic and visibility—have become important spaces for public expression and policy discussion. These accounts reach large audiences and maintain high engagement, making them well-suited for observing shifts in public sentiment. Our aim is to capture both the content posted by these accounts and the real-time public engagement their posts generate.
@@ 2 Data and Measurement
Figure 1: Data Collection The Figure 1(a) illustrates our data collection process. We actively monitored the timelines of 165 government and media accounts. At time $T_{0}$, if a new post appeared, the system began tracking it and its top comments. Snapshots were taken every five minutes for the first 24 hours and every 24 hours for the following 10 days. Figure 1(b) provides an example of the information collected in each snapshot, including the publisher, publication time, post content, and engagement metrics, as well as the top 20 comments and their associated IP location tags. Figure 1(c) shows the number of new posts collected per day from April 18 to May 9, 2022. The vertical red line marks noon on April 28, when the user location disclosure policy was implemented. Figure 1(d) displays the daily distribution of new comments over the same period. Both figures show regular daily fluctuations, indicating consistent user activity. Figures 1(e) and 1(f) report average likes and replies by comment rank. Engagement drops sharply after the first page, and comments on following pages receive minimal interaction. Although we also estimate the causal effect on the total number of comments, our primary focus is on first-page comments, which account for the majority of meaningful engagement.
2.1 High-frequency Weibo Data

### v3

@@ Abstract
We examine the behavioral impact of a user location disclosure policy on Sina Weibo, China’s largest microblogging platform, using a unique high-frequency dataset of uncensored engagement—including tens of thousands of comments and replies—on prominent government and media accounts. The policy, publicly justified as a measure to curb misinformation and counter foreign influence, was abruptly rolled out on April 28, 2022. Using an interrupted time series design, we find no decline in participation by overseas users. Instead, it significantly reduced domestic engagement with local issues outside users’ home provinces, particularly among critical comments. Evidence indicates this decline was not driven by generalized fear or concerns about credibility, but by a surge in regionally discriminatory replies that raised the social cost of cross-provincial engagement. Our findings suggest that identity disclosure tools can have unintended consequences by activating existing social divisions in ways that reinforce state control without direct censorship.
Keywords Location Disclosure  $\cdot$ Online Behavior  $\cdot$ Social Divisions  $\cdot$ Information Control  $\cdot$ Authoritarianism
@@ 1 Introduction
In October 2021, the Cyberspace Administration of China released draft regulations requiring platforms to display users’ IP-based locations, by province for domestic users and by country for overseas users. Officials framed the policy as a safeguard against misinformation and foreign interference, but critics viewed it as another step to constrain online speech (Reuters, 2022; The New York Times, 2022). On March 17, 2022, less than a month after the Russian invasion of Ukraine, Sina Weibo, China’s largest microblogging site, began tagging the locations of users posting about the war. On April 28, it abruptly extended the policy to all posts and comments. Other major platforms soon followed, and by the end of 2022, location disclosure had become a standard feature across Chinese social media.
@@
This paper examines the behavioral consequences of Weibo’s location disclosure policy. Authoritarian regimes have long relied on censorship (King et al., 2013a; Lorentzen, 2014; Roberts, 2018), surveillance (Esarey and Xiao, 2011; Pan et al., 2023), and orchestrated engagement (Chen and Xu, 2017a, b; Qin et al., 2017). Increasingly, they also deploy identity disclosure tools, such as real-name registration or IP-based location tags, to shape behavior without direct repression. By making identity attributes visible, these tools can alter participation through reputational and social pressures. Prior work shows that regional cues can heighten antagonism and incivility (Peng, 2021; Chen et al., 2021; Yu and Margolin, 2024; Guo et al., 2023), suggesting that disclosure may not only chill expression but also activate existing regional divisions. Yet the causal effects of such measures remain poorly understood.
@@
Figure 1: Data Collection Figure 1(a) illustrates our data collection process. We actively monitored the timelines of 165 government and media accounts. At time $T_{0}$, if a new post appeared, the system began tracking it and its top comments. Snapshots were taken every five minutes for the first 24 hours and every 24 hours for the following 10 days. Figure 1(b) provides an example of the information collected in each snapshot, including the publisher, publication time, post content, and engagement metrics, as well as the top 20 comments and their associated IP location tags. Figure 1(c) shows the number of new posts collected per day from April 18 to May 9, 2022. The vertical red line marks noon on April 28, when the user location disclosure policy was implemented. Figure 1(d) displays the daily distribution of new comments over the same period. Both figures show regular daily fluctuations, indicating consistent user activity. Figures 1(e) and 1(f) report average likes and replies by comment rank. Engagement drops sharply after the first page, and comments on following pages receive minimal interaction. Although we also estimate the causal effect on the total number of comments, our primary focus is on first-page comments, which account for the majority of meaningful engagement.
Studying these effects poses serious empirical challenges. In tightly censored environments like Weibo, observed behavior often reflects censorship and deletion rather than genuine user response. Sensitive posts are removed quickly—3% within 30 minutes and 90% within 24 hours in some cases (King et al., 2013b)—creating severe survivorship bias. Public datasets usually rely on post-censorship crawling, which overrepresents high-follower accounts and relies on narrow keyword lists. Causal inference is also difficult: randomized disclosure is ethically problematic, while observational studies based on deletion flags or self-reported geotags are vulnerable to selection bias.
We address these limitations with a unique high-frequency dataset that continuously monitored 165 prominent government and media accounts and their comment sections at approximately five-minute intervals. This approach captured tens of thousands of comments and replies before censorship or removal, substantially reducing survivorship bias and preserving authentic user responses. A rare Weibo implementation glitch further enabled us to recover pre-treatment geographic identities. Exploiting the abrupt and unannounced expansion of the policy as a natural intervention, we apply an interrupted time series design to estimate its causal effect. By comparing engagement immediately before and after the cutoff under plausible assumptions of no anticipation and smooth potential trends, we provide credible evidence on how identity disclosure reshaped online behavior. Figure 1 illustrates our data collection process, which we will explain in detail in the Methods section.
We find that, contrary to the government’s stated aim, overseas users did not reduce their participation; if anything, they briefly increased their comments on international issues in a reactive backlash. By contrast, domestic users sharply curtailed their involvement in discussions of local issues outside their own provinces. This decline was not explained by shifts in post content but by the withdrawal of out-of-province participants, particularly those offering critical remarks. The new visibility of provincial identity increased the risk of backlash, deepened regional antagonism, and raised the social costs of cross-provincial engagement, thereby narrowing the scope of online discourse without direct censorship.
Beyond the authoritarian context, these findings contribute to a broader literature on how platform architecture shapes online behavior. Studies of credibility labels, state-affiliation tags, and feed-ranking algorithms show that design choices affect exposure, trust, and sharing even without changing beliefs (Flaxman et al., 2016; Guess et al., 2023; Liang et al., 2022; Papakyriakopoulos et al., 2022; Bradshaw et al., 2024). We extend this work by demonstrating that simple identity disclosure can silence participation by activating social cleavages, underscoring the broader importance of platform design in structuring the boundaries of online discourse in both authoritarian and democratic settings.
2 Findings
We now turn to the empirical findings. First, we examine whether the policy achieved its intended goal of deterring participation by overseas users. We then assess its effects on domestic engagement, distinguishing between local and non-local issues, and explore the mechanisms driving these patterns.
The policy was designed to deter participation by overseas users with “malicious intent,” who were believed to interfere in domestic discourse by spreading misinformation and expressing views that challenge the official narrative. The expectation was twofold: that public location tags would help domestic audiences judge the credibility of online content, and that overseas users might withdraw out of concern about being discredited or attacked as part of a “malicious foreign force” (see Supplementary Information for background). To evaluate these claims, we distinguish between posts on international affairs, defined as those mentioning at least one foreign location, and non-international posts, which cover domestic issues or general topics without foreign reference.
(a) #Comments by Oversea Users
(b) #Comments by Domestic Users
(c) #Comments by Domestic Users
on Non-International issues
(d) Proportion of Local Posts Covering Positive & Negative Events
(e) #Comments by Domestic Users: by User Origin
(f) #Critical Comments by Domestic Users: by User Origin
Figure 2: Engagement before and after User Location Disclosure. Figures 2(a) and 2(b) show smoothed trends in the number of first-page comments from overseas and domestic users, respectively, on posts about international and non-international topics. Comment counts are aggregated over the first 48 hours after posting, and the vertical red line marks the implementation of the IP location disclosure policy at noon on April 28, 2022. Figure 2(a) shows no decline in overseas engagement. Comments on international topics rose temporarily before returning to baseline, while comments on non-international topics remained stable. By contrast, Figure 2(b) shows a sharp and sustained drop in domestic comments on non-international posts, with no change on international topics. Figure 2(c) confirms that this decline is driven by posts about local issues—defined as those mentioning a province name—while engagement with non-local topics remains stable. Figure 2(d) shows no significant change in the distribution of local post content by sentiment, ruling out shifts in post type as an explanation. Figures 2(e) and 2(f) further show that the drop in engagement is driven primarily by out-of-province users and is accompanied by a reduction in dissenting comments. In-province engagement also declined but rebounded more quickly. These patterns suggest that user location disclosure discouraged nonlocal participation in local discussions. Confidence intervals are constructed based on cluster-robust standard errors clustered at the account level.
Figure 2(a) tests this deterrence hypothesis directly. The blue and red smoothed lines show the average number of comments from overseas users (based on platform-identified IP locations) on international and non-international posts, respectively, before and after the policy took effect. Contrary to expectations, there is no evidence of reduced overseas engagement. Comments on international topics rose sharply, from an average of 0.41 to 0.78 per post, before gradually returning to baseline. This short-lived surge suggests backlash rather than retreat. This result is unlikely to be driven by VPN masking because 99.1% of user IP tags remained unchanged across the policy shift, and analysis in Section G of the Supplementary Information shows that overseas users continued to comment on sensitive topics such as the Russian invasion of Ukraine, despite increased attacks from domestic commenters. On non-international posts, overseas engagement remained stable, further weakening the deterrence hypothesis. Taken together, these results indicate that the policy failed in its stated goal of discouraging overseas participation. Note that the red line remains consistently lower than the blue line, but has much narrower confidence intervals. This is because overseas users are more likely to comment on posts related to international affairs; however, such posts are less frequently made by propaganda accounts compared to non-international posts, which explains the wider confidence intervals around the blue line.
Turning to domestic users, Figures 2(b) shows a sharp and sustained decline in engagement. As in Figure 2(a), the blue and red smoothed lines represent the average number of comments on international and non-international posts, respectively—this time from domestic users, based on their platform-identified user locations. We see that comment volumes on international posts remained stable, but those on non-international posts fell immediately by about 50%. Again, the wider confidence intervals around the blue line reflect the fact that government and media accounts post far fewer international than non-international topics, resulting in greater variability in the estimated trend. Disaggregating further, in Figure 2(c), we find that this drop is concentrated in discussions of local issues, defined as posts mentioning a Chinese province. Engagement with posts concerning local affairs plummeted from an average of 8.21 to 0.42 comments per post after the policy (purple lines), while participation on non-local topics remained stable (red lines). [^1] Consistent with this finding, Figure G26 in the Supplementary Information shows that aggregate engagement metrics, including the total number of comments, likes, and reports, on posts about local issues also declined following the policy. However, using the aggregate metrics, we cannot distinguish whether this decrease is driven by domestic or international users.
In sum, while the policy did not deter overseas users, it sharply curtailed domestic engagement with local issues outside users’ home provinces. The intended targets were unaffected, but curiously, domestic discourse contracted in precisely the spaces where local identity cues were made most salient.
2.2 Decline in discussions on local affairs were driven by out-of-province users
Why did user engagement on posts about local issues fall so sharply after the policy took effect? We first test whether the decline reflects changes in what government and media accounts were posting. If these accounts shifted toward fewer negative or politically sensitive stories, reduced engagement could follow. Yet Figure 2(d) shows that the distribution of local posts across positive, neutral, and negative events remained stable before and after implementation. These negative events include major accidents, scandals, and COVID-related lockdowns. Moreover, Figure D5 in the Supplementary Information demonstrates that engagement fell for both positive and negative events alike. This rules out changes in post content as the main explanation.
If not content, then what may be the drivers? We consider two possibilities: a chilling effect, where users feared state repression once their locations became visible, or heightened social pressure, where visible location tags altered the dynamics of local discussions. To probe these possibilities, we compare “in-province” users—those who comment on posts about their own province—with “out-of-province” users—those who comment on posts about other regions.
Figure 2(e) shows that the steepest decline came from out-of-province users. Their average comments per post fell from 6.10 to 0.42 and remained depressed in the days that followed. In-province users also reduced participation, but the drop was smaller (from 1.85 to 0.15 comments per post) and engagement rebounded quickly. This pattern indicates that the contraction of local discussions was driven mainly by the withdrawal of outsiders. The composition of comments also shifted. Figure 2(f) reveals that the share of critical or dissenting comments on local posts declined significantly once IP locations were displayed, largely because out-of-province users—who had previously contributed disproportionately to criticism—stopped commenting.
These findings make a chilling effect imposed by the state an unlikely primary explanation. If fear of repression were the main driver, we would expect users to be most cautious when commenting on affairs in their own province, where risks of being identified and penalized are highest. Instead, we observe the opposite pattern: out-of-province users withdrew, while in-province users continued to participate. Consistent with this interpretation, domestic users did not reduce their engagement on international or national (non-local) issues, including posts that drew critical comments (Figures 2(b) and 2(c)).
In sum, the policy did little to deter overseas users but substantially dampened domestic discussion of local issues. The sharpest declines came from out-of-province users, who, once their locations became visible, appeared less willing to criticize regions where they did not reside. These results suggest that social pressure rooted in regional divisions, rather than fear of repression, may have played a central role. We turn next to the role of regional identity in shaping online discourse.
Why did domestic users—especially those commenting on other provinces—withdraw from local discussions after IP location disclosure? Beyond the chilling effect already considered, we evaluate two additional mechanisms.
The first possibility is concerns of loss of credibility. Location tags might have signaled that out-of-province commenters lacked local knowledge, reducing the perceived legitimacy of their contributions. This interpretation aligns with the policy’s stated rationale of curbing misinformation. If credibility concerns were the driving factor, however, we would expect similar declines among overseas users commenting on domestic issues and domestic users commenting on international issues. Yet neither group reduced engagement (Figure 2(a), red line; Figure 2(b), blue line). Therefore, credibility alone cannot explain the observed pattern.
(a) Intensified Regional Discrimination
(b) Less Comments with Reply
(c) Reduced Coefficient of Variation
A more plausible explanation is fear of backlash from other users. Location tags made provincial origins visible, heightening the salience of regional identity and turning comment sections into sites of intergroup tension. Before the policy, criticism of local affairs often appeared anonymous or broadly directed. Afterward, even neutral observations could be read as attacks from outsiders. The case study in Section H of the Supplementary Information illustrates this dynamic: a Fujian user commenting on population decline in the northeast invoked stereotypes about “mafia” in Heilongjiang, Liaoning, and Jilin, prompting retaliatory replies about fraud and scams in Fujian. Location tags thus transformed casual remarks into interprovincial confrontations.
To assess whether such exchanges were systemic, we applied an LLM to detect regionally discriminatory language in second-tier replies. Figure 3(a) shows a sharp increase in discriminatory replies after the rollout, driven almost entirely by cross-provincial interactions (Figure G25, SI). Within-province exchanges remained stable. These results indicate that the policy deepened regional divisions and raised the reputational costs of participating in discussions about other regions.
This dynamic helps explain the broader decline in comment activity on local posts. As regional identity became more visible, users faced greater risks of confrontation—not just for critical remarks but for engaging at all. Even neutral comments were more likely to attract hostile replies, discouraging participation and reducing the likelihood of responses. Figure 3(b) documents this shift: after disclosure, a smaller share of comments attracted replies, signaling thinner, more cautious engagement.
We further measure this contraction in discourse by examining the rotation of first-page comments using the coefficient of variation (CoV) of comment floor numbers, a proxy for dynamism in the comment section. Figure 3(c) shows a clear post-policy drop in CoV. With users more hesitant to comment—especially across provincial boundaries—top comments were displaced less often, producing a more static and less interactive hierarchy, hence the drop in the CoV.
In sum, location disclosure not only reduced overall engagement but also reshaped its structure. By making regional identity salient, it elevated the risks of cross-provincial participation, curtailed dissent, and produced a narrower, more fragmented conversation.
3 Discussion
Our study examined the impact of Weibo’s location disclosure policy, focusing on how the abrupt introduction of IP-based location tags reshaped online engagement. The policy, framed as a tool to curb misinformation and foreign influence, did little to deter overseas users. Instead, it curtailed domestic participation in discussions of local issues, especially critical engagement by out-of-province users. Evidence suggests that this decline was not driven by fear of repression but by heightened exposure to regionally discriminatory backlash once users’ provincial origins became visible.
These findings highlight how authoritarian regimes can shape discourse without direct censorship. By embedding identity disclosure into platform design, authorities enable peer-based sanctions that suppress participation and narrow the scope of expression. Which such mechanisms may stabilize discourse in the short run, they risk muting early signals of discontent, with potential costs for the quality of governance (Chen and Xu, 2017b).
Several limitations qualify our results. First, the analysis focuses on highly visible comment sections of government and media accounts, which may not reflect how users behave in smaller or less monitored spaces. Second, although our high-frequency collection reduces survivorship bias, we did not capture other forms of engagement such as reposting. In addition, IP tags provide only province- or country-level origins, limiting our ability to study how other identity dimensions—such as gender, class, or ethnicity—intersect with location in shaping discourse.
The scope of inference is also bounded by the sudden, unannounced rollout of the policy in an authoritarian setting. The causal effects we are able to identify are immediate behavioral shifts; longer-term adaptations may differ as users adjust or migrate to alternative venues. Future work could explore whether similar dynamics emerge in democratic contexts, where censorship risks are lower but identity cues may still intensify polarization, and extend analysis to other forms of identity disclosure beyond geography.
Nevertheless, our findings contribute to a growing literature on platform architecture and user behavior. Prior research shows that design features such as algorithmic curation, identity labeling, and credibility signals affect what users see, share, and trust (Guess et al., 2023; Liang et al., 2022; Nassetta and Gross, 2020; Papakyriakopoulos et al., 2022; Bradshaw et al., 2024). We add to this work by showing that identity cues alone can alter participation even without removing content or correcting information. This underscores how design decisions—whether in democracies or authoritarian states—shape the boundaries of public discourse in subtle but consequential ways.
4 Methods
This study draws on a high-frequency dataset and a quasi-experimental design to assess the causal impact of Weibo’s abrupt rollout of IP-based location tags. The main challenges are capturing genuine user behavior in a tightly censored environment and separating the effect of the policy from other influences. We address these by constructing a real-time monitoring system that tracked posts and comments across the policy change and by applying an interrupted time series design that exploits the unanticipated timing of implementation.
4.1 Data and Measurement
We collaborated with industry experts to design a real-time data collection system that monitored 165 prominent Weibo accounts affiliated with the Chinese government. These include official accounts of government offices and state media outlets—such as newspapers, television stations, and major media outlets—identified in a series of reports on institutional influence in 2021.222These reports include multiple editions of Government Weibo Influence Reports (政务微博影响力报告), Sina Government Weibo Reports (新浪政务微博报告), and Weibo Trending Topics Data Reports (微博热搜榜数据报告). See Section B in the Supplementary Information for a full list of accounts we collected data from. We focused on these accounts because, following the government’s 2013 crackdown on key opinion leaders (Pan et al., 2025), the comment sections of these accounts—due to high traffic and visibility—have become important spaces for public expression and policy discussion. These accounts reach large audiences and maintain high engagement, making them well-suited for observing shifts in public sentiment. Our aim is to capture both the content posted by these accounts and the real-time public engagement their posts generate.
High-frequency Weibo Data